OpenStudy (anonymous):

Which of the following is the solution to the equation 4 over 5n = 20? n = 4 n = 12 n = 16 n = 25

OpenStudy (anonymous):

n =25

OpenStudy (kendricklamar2014):

Here are the steps @mehdimunk \[\frac{ 4 }{ 5n } = 20 \] Multiplu both sides by 5n\[4=100n\]

OpenStudy (kendricklamar2014):

\[4=100n\] Fllp equation \[100n=4\]

OpenStudy (kendricklamar2014):

Divide both sides by 100 \[\frac{ 100n }{ 100 } = \frac{ 4 }{ 100 }\] = \[n = \frac{ 1 }{ 25 }\]

OpenStudy (kendricklamar2014):

\[n = \frac{ 1 }{ 25 } ~ is~the~same~as~ n = 25\]

OpenStudy (kendricklamar2014):

Answer:\[n = 25\]

OpenStudy (anonymous):

my answer is the right one

OpenStudy (kendricklamar2014):

Yes, but you gave a direct answer.

OpenStudy (kendricklamar2014):

@yajna

OpenStudy (anonymous):

what i did was just replacing the n values in 4/5n =20

OpenStudy (kendricklamar2014):

Still, you gave a direct answer. Next time, explain first!

OpenStudy (anonymous):

no need to be rude dude

OpenStudy (anonymous):

Kendrick, you're steps are wrong though. 1/25 doesn't equal 25

OpenStudy (anonymous):

(4/5)n=20

OpenStudy (anonymous):

n=20(5/4)

OpenStudy (anonymous):

n=25